# Parameter-Efficient Fine-Tuning in Large Models: A Survey of Methodologies

### v2

@@ III-C1 Unstructural Masking
where $\hat{F}_{\theta}$ represents the empirical Fisher information, $n$ is the number of most regressing training examples, $p_{\theta}(y_{i}|x_{i})$ is the output probability for the given input $x_{i}$ and parameters $\theta$, and the sum is taken over the top $n$ regressing examples as determined by their loss $L_{tr}$ during training. LT-SFT [159] introduces a composable sparse fine-tuning method for cross-lingual transfer learning. It learns sparse, real-valued masks based on a variant of the Lottery Ticket Hypothesis (LTH). Task-specific masks are derived from supervised data in the source language, while language-specific masks are obtained through masked language modeling in the target language. These masks are composed with the pre-trained model to enable zero-shot cross-lingual transfer. The sparsity of the masks reduces parameter overlap and interference, improving modularity and preventing overfitting. SAM (Second-order Approximation Method) [160] approximates the original optimization problem using a second-order Taylor expansion to make it analytically solvable, and directly determines the parameters to optimize by solving the approximation function, which is formulated as:
@@
subject to $\|M\|_{0}=\lfloor mp\rfloor;M_{ij}=0,\forall i\neq j;M_{ii}\in\{0,1\}$, where $\theta_{0}$ are the pre-trained parameters, $\Delta\theta$ is the difference vector, $M$ is the parameter mask matrix, $L$ is the loss function, $\nabla L(\theta_{0})$ is the gradient of the loss function at $\theta_{0}$, and $H$ is an approximated diagonal Hessian matrix. Child-tuning [161] updates only a subset of parameters, referred to as the child network, during fine-tuning while masking out the gradients of the remaining parameters in the backward pass, which can be formulated as:
@@
where $w_{t}$ represents the model parameters at the $t^{\text{th}}$ iteration, $\eta$ is the learning rate, $L(w_{t})$ is the loss function, and $M_{t}$ is a 0-1 mask indicating the child network. U-MAM [155] is an unstructured neural architecture search approach for parameter-efficient tuning of large pre-trained language models. It involves pruning a dense low-rank update from an initial parameter-efficient tuning architecture to find an efficient subset of parameters to fine-tune. Threshold-Mask [162] learns selective binary masks for pre-trained language model weights without fine-tuning, where each linear layer $W_{l}$ is associated with a real-valued matrix $M_{l}$ initialized randomly, and a binary mask $M_{l}^{bin}$ is obtained by applying a thresholding function, used to select important weights: $(m_{l}^{bin})_{i,j}=1(m_{l,i,j}\geq\tau)$ with $m_{l,i,j}\in M_{l}$ and the global thresholding hyperparamter $\tau$, and the masked weights are computed as $\hat{W}_{l}=W_{l}\odot M_{l}^{bin}$, with $M_{l}$ updated during training via the straight-through estimator: $M_{l}\leftarrow M_{l}-\eta\frac{\partial L(\hat{W}_{l})}{\partial M_{l}^{bin}}$. LoRAPrune [163] approximates the importance of each parameter in the pre-trained model weights $W_{0}$ by utilizing the gradients of the low-rank matrices $A$ and $B$, which are then used to perform structured pruning in an iterative and progressive manner, efficiently reducing the model’s size while maintaining performance.
@@ III-D Hybrid PEFT
Due to the significant performance differences of different types of PEFT methods on various tasks, many studies aim to enhance model performance by combining the advantages of different types of PEFT methods. These research efforts are summarized as Hybrid PEFT methods.
@@ III-F Multi-task PEFT
The previously introduced PEFT methods were mainly designed for single downstream task. This section focuses on PEFT for multi-task learning.
@@ IV Applications of PEFT
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=6.0em,font=, where level=2text width=5.5em,font=, where level=3text width=5.5em,font=, [ PEFT Methods for Other PLMs, ver [ Vision Models [ Image
Classification [ VP [197], VPT [198], NOAH [170], Convpass [199], AdaptFormer [200], DAM-VP [201], ILM-VP [202],
EVP [203], VQT [204], FacT [205], DTL [206], LION [207] , leaf, text width=34.1em ] ] [ Dense
Prediction [ Polyhistor [208], ViT-Adapter [209], SAN [210], LoRand [211] , leaf, text width=34.1em ] ] ] [ Diffusion Models [ Generation by
@@
Fine-tuning [ DreamBooth [212], Textual Inversion [213], DreamArtist [214], Extended Textual Inversion [215],
DiffFit [216], Cones [217], SVDiff [218], LyCORIS [219], DiffuseKronA [220], OFT [221] , leaf, text width=34.1em ] ] [ Controllable
Generation [ Sketch-guided Diffusion [222], ControlNet [223], T2I-Adapter [224], Uni-ControlNet [225],
IP-Adapter [226] , leaf, text width=34.1em ] ] ] [ MLLM [ BLIP-2 [227], LLaVA [228, 229], Flamingo [230], LLaMA Adapter [231], CogVLM [232], Q-Former [227, 233], GPT-4 [30] , leaf, text width=41.2em ] ] ]
Figure 8: Taxonomy of PEFT Methods for Other Types of Foundation Models
@@ IV-A2 Dense Prediction
Dense prediction, encompassing tasks such as image segmentation, object detection, depth estimation, etc., is another crucial task in the field of 2D vision. Unlike image classification tasks, which typically generate a single prediction label for an entire image, dense prediction tasks require making predictions for every pixel in the image, usually resulting in an output image with the same resolution as the input image. Fine-tuning pre-trained models from image classification is a common approach for dense prediction tasks. With the application of PEFT methods in vision tasks, various PEFT methods tailored for dense prediction tasks have been proposed.
@@ IV-B2 Controllable Generation
Controllable generation primarily involves adding control sources beyond the prompt to guide the image generation. These control sources can include sketches, keypoints, or other forms of guidance to shape the generated output more precisely.
@@
$\text{F}_{\text{DN}r}(\text{Z}_{r},c_{l})=\text{norm}(\text{Z}_{r})\cdot(1+\text{conv}_{\gamma}(\text{zero}(h_{r}(c_{l}))))\\ +\text{conv}_{\beta}(\text{zero}(h_{r}(c_{l})))\enskip,$ (35)
@@ IV-C PEFT in MLLM
The parameter-efficient fine-tuning of MLLM primarily focuses on the model connector. It is because maintain consistency for both multimodal and textual data is challenging. As a consequence, a modal connector is serially connected right before the LLM, converting multimodal embeddings into understandable text prompt tokens for the LLM. Training the model connector on peft dataset bridges the gap between different modal data while ensuring consistency in the input to the LLM.
@@
Figure 9: Modal Connector Design

### v3

@@ III PEFT Taxonomy
Category Main Idea Representative Methods #Trainable Parameters Applications Advantages Limitations
Additive Add trainable components, freeze original. Sequential Adapter [15], Prefix-tuning [16], $(IA)^{3}$ [97], IPA [165] #Params of additional modules Single-task, rapid adaptation. Minimal updates, flexible insertion. Computational overhead, design-sensitive.
Reparameterized Low-rank decomposition, tune low-rank matrices. LoRA [19], AdaLoRA [20], DoRA [175] #Params of low-rank matrices Large-scale, efficient updates. Fewer parameters, no inference latency. Low-rank constraints, hyperparameter tuning.
Selective Update subsets (e.g., biases, masked params). U-Bitfit [184], FAR [193] #Params of selected subsets (e.g., biases, masked params) Resource-constrained environments. Critical updates, low memory. Task-sensitive, parameter selection.
Hybrid Combine multiple PEFT methods dynamically. UniPELT [196], MAM-Adapter [144] #Params of used PEFT modules Complex tasks, multimodal. Task flexibility, improved performance. High complexity, search overhead.
Quantization Quantize model, enable efficient tuning. QLoRA [205], BitDelta [210] #Params of used PEFT modules Edge devices, low resources. Low storage, low-precision inference. Precision loss, quantization balance.
Multi-task Share parameters and dynamic adapters for multi-task. AdapterFusion [211], SPoT [217], MOELoRA [223] #Params of shared and task-specific modules Multi-task, cross-task knowledge. Redundant reduction, task transfer. Task conflicts, routing complexity.
TABLE V: An Overview of Different Types of PEFT Methods: Main Idea, Number of Trainable Parameters, Applications, and Limitations.
Model PEFT Type PEFT Method #TPs CoLA SST2 MRPC STS-B QQP MNLI QNLI RTE
RoBERTa-base FT 124.6M 59.07 92.89 88.24/91.58 90.87/90.61 90.81/87.72 86.27 91.07 72.2
Additive AdapterS 7.41M 63.32 94.31 90.44/93.18 91.25/90.94 90.81/86.55 87.33 92.06 73.56
Prefix-tuning 0.96M 59.31 93.81 87.25/91.03 88.48/88.32 87.75/84.09 85.21 90.77 54.51
(IA)3 0.66M 59.58 93.92 87.00/90.52 90.30/90.32 87.99/84.10 83.95 90.88 71.12
Reparameterized LoRA 0.89M 62.09 94.04 87.50/90.68 90.66/90.83 88.83/85.21 86.54 92.02 72.92
AdaLoRA 1.03M 59.82 93.92 87.99/91.33 90.83/90.73 88.58/84.98 86.26 91.43 70.04
Selective BitFit 0.69M 61.32 94.72 89.22/92.41 90.34/90.27 88.12/84.11 84.64 91.09 77.98
Child-Tuning / 60.33 93.58 89.22/92.20 91.14/90.93 90.98/88.04 87.4 92.2 77.62
Hybrid MAM Adapter 46.78M 61.42 94.87 89.31/92.21 90.74/90.42 88.31/83.20 86.63 90.19 72.62
RoBERTa-large FT 355.3M 65.78 95.54 89.22/92.28 91.74/91.76 89.30/86.68 89.42 93.61 81.23
Additive AdapterS 19.77M 67.03 96.37 89.94/92.54 92.58/92.42 92.19/88.50 91 94.31 85.25
Prefix-tuning 2.03M 59.01 95.76 88.24/91.37 90.92/91.07 88.88/85.45 89.3 93.32 74.01
(IA)3 1.22M 61.15 94.61 86.52/90.33 92.22/92.03 89.45/86.25 88.63 94.25 81.23
Reparameterized LoRA 1.84M 64.47 96.67 87.50/91.19 91.66/91.44 90.15/86.91 90.76 95 79.78
AdaLoRA 2.23M 65.85 94.95 89.46/92.34 92.05/91.80 89.60/86.30 90.36 94.62 77.98
Selective BitFit 1.32M 68.01 96.1 90.93/93.38 91.93/91.77 89.48/86.43 89.98 94.47 87.73
Child-Tuning / 63.08 95.07 90.69/93.43 92.36/92.18 91.52/88.75 35.45 93.15 86.25
Hybrid MAM Adapter 122.2M 67.39 95.81 90.12/92.77 92.44/92.18 90.87/86.65 90.62 94.31 86.62
DeBERTaV3-base FT / 69.2 95.3 89.5/93.3 91.6/91.1 92.4/89.8 90.5 94 82
Quantization QLoRA / N.A. 86.5 73.8/82.8 83.0/82.8 86.8/82.3 75.4 82.4 55.9
LoftQ / 37.4 90.2 83.8/88.6 87.1/86.9 90.3/86.9 84.7 86.6 61.4
TABLE VI: Performance evaluation across various PEFT methods for fine-tuning common base models (RoBERTa-base, RoBERTa-large, and DeBERTaV3-base) on the GLUE benchmark. All performance metrics are cited from prior published works [25, 209]. Metrics may vary by task: Matthews correlation for COLA, accuracy/F1 score for MRPC and QQP, Pearson/Spearman correlation for STS-B, average matched accuracy for MNLI, and accuracy for the remaining tasks. Higher metric values indicate superior performance. #$\#TP$ denotes the number of trainable parameters for each method.
@@ III-C1 Unstructural Masking
Figure 8: Illustration of three representative unstructural masking methods
@@
where $\hat{F}_{\theta}$ represents the empirical Fisher information, $n$ is the number of most regressing training examples, $p_{\theta}(y_{i}|x_{i})$ is the output probability for the given input $x_{i}$ and parameters $\theta$, and the sum is taken over the top $n$ regressing examples as determined by their loss $L_{tr}$ during training. LT-SFT (see Figure 8 (c)) [188] introduces a composable sparse fine-tuning method for cross-lingual transfer learning. It learns sparse, real-valued masks based on a variant of the Lottery Ticket Hypothesis (LTH). Task-specific masks are derived from supervised data in the source language, while language-specific masks are obtained through masked language modeling in the target language. These masks are composed with the pre-trained model to enable zero-shot cross-lingual transfer. The sparsity of the masks reduces parameter overlap and interference, improving modularity and preventing overfitting. SAM (Second-order Approximation Method) [189] approximates the original optimization problem using a second-order Taylor expansion to make it analytically solvable, and directly determines the parameters to optimize by solving the approximation function, which is formulated as:
@@
subject to $\|M\|_{0}=\lfloor mp\rfloor;M_{ij}=0,\forall i\neq j;M_{ii}\in\{0,1\}$, where $\theta_{0}$ are the pre-trained parameters, $\Delta\theta$ is the difference vector, $M$ is the parameter mask matrix, $L$ is the loss function, $\nabla L(\theta_{0})$ is the gradient of the loss function at $\theta_{0}$, and $H$ is an approximated diagonal Hessian matrix. Child-tuning (see Figure 8 (b)) [190] updates only a subset of parameters, referred to as the child network, during fine-tuning while masking out the gradients of the remaining parameters in the backward pass, which can be formulated as:
@@
where $w_{t}$ represents the model parameters at the $t^{\text{th}}$ iteration, $\eta$ is the learning rate, $L(w_{t})$ is the loss function, and $M_{t}$ is a 0-1 mask indicating the child network. U-MAM [184] is an unstructured neural architecture search approach for parameter-efficient tuning of large pre-trained language models. It involves pruning a dense low-rank update from an initial parameter-efficient tuning architecture to find an efficient subset of parameters to fine-tune. Threshold-Mask [191] learns selective binary masks for pre-trained language model weights without fine-tuning, where each linear layer $W_{l}$ is associated with a real-valued matrix $M_{l}$ initialized randomly, and a binary mask $M_{l}^{bin}$ is obtained by applying a thresholding function, used to select important weights: $(m_{l}^{bin})_{i,j}=1(m_{l,i,j}\geq\tau)$ with $m_{l,i,j}\in M_{l}$ and the global thresholding hyperparamter $\tau$, and the masked weights are computed as $\hat{W}_{l}=W_{l}\odot M_{l}^{bin}$, with $M_{l}$ updated during training via the straight-through estimator: $M_{l}\leftarrow M_{l}-\eta\frac{\partial L(\hat{W}_{l})}{\partial M_{l}^{bin}}$. LoRAPrune (see Figure 8 (a)) [192] approximates the importance of each parameter in the pre-trained model weights $W_{0}$ by utilizing the gradients of the low-rank matrices $A$ and $B$, which are then used to perform structured pruning in an iterative and progressive manner, efficiently reducing the model’s size while maintaining performance.
@@ III-D Hybrid PEFT
Due to the significant performance differences of different types of PEFT methods on various tasks, many studies aim to enhance model performance by combining the advantages of different types of PEFT methods. These research efforts are summarized as Hybrid PEFT methods. A representative hybrid PEFT method, known as MAM-Adapter, is illustrated in Figure 9.
Figure 9: Illustration the principles of MAM-Adapter, a representative hybrid PEFT method
@@ III-E Quantization PEFT
Figure 10: Illustration of the difference among LoRA, QLoRA and QA-LoRA
@@ III-F Multi-task PEFT
The previously introduced PEFT methods were mainly designed for single downstream task. This section focuses on PEFT for multi-task learning. Figure 11 illustrates three multi-task PEFT approaches: AdaMix (Adapter-based), ATTEMPT (Soft Prompt-based), and MOELoRA (LoRA-based).
Figure 11: Illustration of three representative multi-task PEFT methods: AdaMix (Adapter-based), ATTEMPT (Soft Prompt-based), and MOELoRA (LoRA-based)
@@ IV Applications of PEFT
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=7.0em,font=, where level=2text width=6em,font=, where level=3text width=5.5em,font=, [ PEFT Methods for Other PLMs, ver [ Vision Models [ Image
Classification [ VP [226], VPT [227], NOAH [199],
Convpass [228], AdaptFormer [229],
DAM-VP [230], ILM-VP [231], EVP [232],
VQT [233], FacT [234], DTL [235],
LION [236] , leaf, text width=34.1em ] ] [ Dense
Prediction [ Polyhistor [237], ViT-Adapter [238], SAN [239],
LoRand [240] , leaf, text width=34.1em ] ] ] [ Diffusion Models [ Generation by
@@
Fine-tuning [ DreamBooth [241], Textual Inversion [242],
DreamArtist [243], Extended Textual Inversion [244],
DiffFit [245], Cones [246], SVDiff [247],
LyCORIS [248], DiffuseKronA [249], OFT [250] , leaf, text width=34.1em ] ] [ Controllable
Generation [ Sketch-guided Diffusion [251], ControlNet [252],
T2I-Adapter [253], Uni-ControlNet [254],
IP-Adapter [255] , leaf, text width=34.1em ] ] ] [ MLLM [ BLIP-2 [256], LLaVA [257, 258], Flamingo [259],
LLaMA Adapter [260], CogVLM [261],
Q-Former [256, 262], GPT-4 [30] , leaf, text width=41.7em ] ] ]
Figure 12: Taxonomy of PEFT Methods for Vision Models, Diffusion Models and MLLM
This section presents a comprehensive overview of PEFT methodologies specifically developed for several prominent applications, categorized as follows: PEFT in Vision Models (IV-A), which primarily focuses on adapting pretrained vision models to specialized computer vision tasks (e.g., image classification, image segmentation, object detection, and depth estimation); PEFT in Diffusion Models (IV-B), which addresses the adaptation of diffusion models for vision generation tasks; and PEFT in MLLM (IV-C), which emphasizes training model connectors on domain-specific datasets to bridge multimodal data discrepancies while maintaining input consistency for LLMs. For a structured overview of these applications and their corresponding recommended PEFT techniques, refer to Figure 12.
@@ IV-A1 Image Classification
In this subsection, we introduce PEFT methods for image classification tasks in vision models. Figure 13 illustrates the principles of three representative PEFT methods discussed in this subsection.
Figure 13: Illustration of the principles of three PEFT methods for image classification: VPT (Soft Prompt-based), AdapterFormer (Adapter-based), and FacT (LoRA-based). ×$\times_{i}$ in FacT is mode-$i$ product.
@@ IV-A2 Dense Prediction
Dense prediction, encompassing tasks such as image segmentation, object detection, depth estimation, etc., is another crucial task in the field of 2D vision. Unlike image classification tasks, which typically generate a single prediction label for an entire image, dense prediction tasks require making predictions for every pixel in the image, usually resulting in an output image with the same resolution as the input image. Fine-tuning pre-trained models from image classification is a common approach for dense prediction tasks. With the application of PEFT methods in vision tasks, various PEFT methods tailored for dense prediction tasks have been proposed. Figure 14 illustrates a representative PEFT method for dense prediction.
Figure 14: Illustration of a representative PEFT method for dense prediction: Polyhistor
@@ IV-B2 Controllable Generation
Controllable generation primarily involves adding control sources beyond the prompt to guide the image generation. These control sources can include sketches, keypoints, or other forms of guidance to shape the generated output more precisely. A representative implementation of controllable generation method is shown in Figure 15
Figure 15: Illustration of the principle of ControlNet, a representative implementation of a controllable generation method
@@
$\text{F}_{\text{DN}r}(\text{Z}_{r},c_{l})=\text{norm}(\text{Z}_{r})\cdot(1+\text{conv}_{\gamma}(\text{zero}(h_{r}(c_{l}))))+\text{conv}_{\beta}(\text{zero}(h_{r}(c_{l})))\enskip,$ (37)
@@ IV-C PEFT in MLLM
The parameter-efficient fine-tuning of MLLM primarily focuses on the model connector. It is because maintain consistency for both multimodal and textual data is challenging. As a consequence, a modal connector is serially connected right before the LLM, converting multimodal embeddings into understandable text prompt tokens for the LLM. Training the model connector on PEFT dataset bridges the gap between different modal data while ensuring consistency in the input to the LLM. As a representative PEFT approach within the MLLM framework, the schematic diagram of LLaMA-Adapter [260] is illustrated in Figure 16.
Figure 16: Illustration of the principle of LLaMA-Adapter, which is a representative PEFT method in MLLM
@@
Figure 17: Modal Connector Design: This figure shows two different mainstream design of the modal connector in MLLM. The first one, a simple MLP for converting the modal. The second one, layers with cross attention and query tokens for training.
@@
Task Model PEFT Method #TPs (M) Result
CIFAR 100 CIFAR 10 Flowers Food EuroSAT SUN DMLab SVHN Pets DTD RESISC CLEVR
Image Classification CLIP FT 151.28 82.1 95.8 97.4 87.8 99 79 63.5 95.7 88.5 72.3 98.1 94.4
VP 0.07 75.3 94.2 62 83.2 95.6 68.4 41.9 88.4 86.5 57.1 84.1 81.4
VPT 0.064 76.6 95 76.2 84.7 94.6 69.3 48.4 86.1 92.1 61.6 84.3 58.6
EVP 0.062 81.2 96.6 82.3 84.1 97.6 71 62.3 90.5 90 68.4 89.7 75.9
Seg. H.Part Sal. Normals.
Dense Prediction Swin Transformer -Tiny Single-task FT 112.62 67.21 61.93 62.35 17.97
Multi-task FT 30.06 68.71 62.13 64.18 17.35
Bitfit 2.85 68.57 55.99 60.64 19.42
Relative bias 2.64 63.51 52.35 57.74 21.07
VPT-shallow 2.57 62.96 52.27 58.31 20.9
VPT-deep 3.43 64.35 55.24 58.15 21.07
PHM layer 3.14 68.55 56.28 60.35 19.23
Compacter 2.78 68.38 56.69 59.47 19.54
Compacter++ 2.66 67.26 55.69 59.47 19.54
LoRA 2.87 67.26 55.69 59.47 19.54
Adapter 11.24 69.21 57.38 61.28 18.83
Low-rank adapter 2.89 68.31 56.53 60.29 19.36
Shared Adapter 4.74 70.21 59.15 62.29 19.26
Hyperformer 75.32 71.43 60.73 65.54 17.77
Polyhistor 8.96 70.87 59.54 65.47 17.47
Polyhistor-Lite 2.96 70.24 59.12 64.75 17.4
Food SUN DF- 20M Caltech CUB- Bird ArtBench Oxford Flowers Standard Cars Average FID
Generation by Few-shot Finetuning DiT-XL-2 FT 673.8 10.46 7.96 17.26 35.25 5.68 25.31 21.05 9.79 16.59
Adapt-Parallel 4.28 13.67 11.47 22.38 35.76 7.73 38.43 21.24 10.73 20.17
Adapt-Sequential 4.28 11.93 10.68 19.01 34.17 7 35.04 21.36 10.45 18.7
BitFit 0.61 9.17 9.11 17.78 34.21 8.81 24.53 20.31 10.64 16.82
VPT-Deep 2.81 18.47 14.54 32.89 42.78 17.29 40.74 25.59 22.12 26.8
LoRA-R8 1.15 33.75 32.33 120.25 86.05 56.03 80.99 164.13 76.24 81.31
LoRA-R16 2.18 34.34 32.15 121.51 86.51 58.25 80.72 161.68 75.35 81.31
DiffFit 0.83 6.96 8.55 17.35 33.84 5.48 20.87 20.18 9.9 15.39
CLIP-T CLIP-I
Controllable Generation CLIP ViT-L/14 Uni-ControlNet (Global Control) 47 0.51 0.74
T2I-Adapter (Style) 39 0.49 0.65
ControlNet Shuffle 361 0.42 0.62
IP-Adapter 22 0.59 0.83
TABLE VII: Performance of PEFT methods in specific applications. All performance metrics are cited from prior published work [232, 237, 245, 253]. Metrics vary by task: 1. Image Classification: 12 datasets with CLIP. 2. Dense Prediction: 4 datasets with Swim Transformer-Tiny. 3. Generation by Few-shot Finetuning: 9 datasets with DiT-XL-2. 4. Controllable Generation: 2 datasets with CLIP ViT-L/14.
To further quantify the performance of different PEFT methods in specific applications, we present Table VII, which compares various methods based on key metrics such as accuracy and the number of trainable parameters across multiple benchmark tasks. Since existing literature does not provide detailed computational cost analysis, we use the number of trainable parameters as an approximate measure of computational efficiency, serving as a practical proxy for resource consumption across different PEFT methods. As shown in Table VII, compared to full fine-tuning, PEFT methods in specific applications significantly reduce the number of trainable parameters while maintaining competitive performance. These results highlight the advantage of PEFT methods in various applications, where they enable efficient adaptation of large models with lower computational and storage costs while preserving task-specific performance.